OpenStudy (idealist10):

Solve y'=(y^2+2xy)/x^2 explicitly.

OpenStudy (idealist10):

y'=y^2/x^2+(2/x)y y'-(2/x)y=y^2/x^2

OpenStudy (anonymous):

WAFFLES lol im lying

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

Do I multiply x^2 by the whole equation?

OpenStudy (anonymous):

No, I think a substitution should work for this one. \[\begin{align*} y'&=\frac{y^2+2xy}{x^2}\\ y'&=\frac{y^2}{x^2}+\frac{2y}{x} \end{align*}\] Let \(u=\dfrac{y}{x}\), or \(y=ux\), then \(y'=u'x+u\): \[\begin{align*} u'x+u&=u^2+2u\\ u'x&=u^2+u\\ \frac{du}{u^2+u}&=\frac{dx}{x} \end{align*}\] Separable... Partial fractions when integrating, or a trig sub.

OpenStudy (idealist10):

I have never used this kind of substitution. Thanks.

OpenStudy (anonymous):

Here's a whole link on it: http://tutorial.math.lamar.edu/Classes/DE/Substitutions.aspx You're welcome!